Pinnacle Plus declared and paid a cash dividend of $7,800 in the current year. Its comparative financial statements, preparedRequired: 1. Compute the gross profit percentage in the current and previous years. Are the current year results better, or wComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ReqComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ReqComplete this question by entering your answers in the tabs below. Required 1 | Required 2 Required 3 Required 4 Required 5 RComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ReqComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ReqComplete this question by entering your answers in the tabs below. Required 1 Required 2 | Required 3 Required 4 Required 5 RComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ReqComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 Req
Answer #1

1. Gross profit ratio = Gross profit / Revenue

a. Current year = 94,000/ 170,000 = 55.3%

b. Previous year = 75,000/147,000 = 51.0%

Current year gross profit is better than previous year

2. Net profit margin = Net profit/ Revenue

a. Current year = 28,560/ 170,000 = 16.8%

b. Previous year =19,040 /147,000 = 13.0%

Current year net profit margin is better than previous year

3. earnings per share (EPS) = Net profit/ No. Of equity shares

a. Current year = 28,560/9,720 = 2.94

b. Previous year = 19,040 / 9,720 = 1.96

EPS has improved from previous year

4. Return on equity = Net profit/ Share holders equity

Average Share holder's Equity = (opening + Closing equity value)/2

Opening Equity value of previous year = 112,000

Profit earned during the previous yearb= 19,040

Closing Equity for previous year = 131,040

Profit earned during the current year = 28,560

Closing Equity of current year = 159,600

Previous year equity average = 121,520

Current year equity average = 145,320

a.Current year return on equity = 28,560/ 121,520 = 20%

b.Previous year return on equity = 19,040/ 145,320 = 24%

Return on equity is worse than previous year.

5. Fixed assets turnover = Sales revnue / Average fixed assets

Average fixed assets previous year = (122,000 + 117,000)/2 = 119,500

Average fixed assets current year = (117,000 + 107,000)/2 = 112,000

a. Current year fixed assets turnover = 170,000 /112,000 = 1.5

b. Previous year ratio = 147,000 /122,000 = 1.2

Ratio has improved from previous year.

6. Debt to asset ratio = Debt/ total asset

A. Current year ratio = 107,300/ 258,160= 0.42

B. Previous year ratio = 86,900/ 217,000= 0.40

Larger portion

7. Interest earned ratio = Income befor interest and taxes/ Interest expenses

A. Current year ratio = 46,000/ 5,200 = 8.8

B. Previous year ratio = 32,400 /5,200 = 6.2

Ratio is better than previous year

8. P/E ratio = Share price/ earnings per share

A. Current year = 27/2.94 = 9.2

B. previous year = 30/1.96 = 15.3

Decline in ratio means investors has become leads optimistic about growth of company